### v4

@@ Abstract
Keywords:
Personalized LLM Cloud-device Collaboration Inference Acceleration.
@@ 1 Introduction
Figure 1: Three setting for model deployment.
@@
  • $\bullet$
    SPA has been thoroughly tested across multiple datasets. Experiments show that our framework performs well on-devices. Additionally, SPA-LST can achieve results close to the original model with lower transmission latency.
@@ 2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by [14], incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA [15] employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning [19] involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
@@ 2.2 On-device ML optimization
Given that a considerable portion of these methods can only be applied to models with small model, they may not produce substantial optimizations in the context of large models. Computational and storage resources on devices may not be sufficient to accommodate LLMs. Therefore, we propose a framework for hybrid cloud server and on-device sequence-to-sequence model. The framework aims to meet the needs of certain downstream tasks and user characteristics by combining the capabilities of the cloud and edge.
3 Side Plugin Adaption
Given the computational and storage capabilities on the edge, our concerns can be broken down into the following points:
  • How much reduction in computation time and space cost can our model achieve?
  • Does its capability improve compared to the original model under these costs?
  • Whether our network latency matches expectations, and whether the overall latency significantly decreases compared to the original model’s cost.
Figure 2: Cloud-base and on-devices collaboration model: SPA.
These are some fundamental requirements and abilities for a small on-device model. Considering the need to learn user features and context-related information for text generation, we propose adding a classifier after the output of the transformer. We explore two settings for this classifier, comparing them to the original edge adapter’s results to assess downstream task support.
In the first setting, a binary choice approach is used, where the edge model is trained from fine-tuning and the output is generated either the origin LLM or side block results. The classifier selects the side whose output better aligns with the desired features, enabling the classifier to choose that path.
The second setting, called the plugin method, involves using the classifier to decide whether to combine the edge model with the base model stored in the cloud. This benefits by allowing the edge model to learn only the differences in features relevant to the current task. General information is then obtained by utilized the base model’s results, which reduced latency cost on edge devices resources.
@@ 3 Side Plugin Adaption
$h=f(x)+\sigma g(x)$ (1)
$\sigma=\arg\max{\textbf{M}_{c}f(x)}$ (2)
We employed a linear layer of $H\times 2$ as our classifier. The advantage of this design lies in our ability to use the Transformer’s output as the hidden state, which is then softmax to a range between [00, $1$], thereby obtaining a specific distribution for classification choices $\sigma$. Through such a linear mapping, the model can learn whether predicting a certain token is better accomplished by selecting the existing model’s output (typically leveraging previously predicted information like "New + York"), or by employing a side-end incremental model (which relies more on user-specific feature information), thus determining the optimal choice. Considering $\sigma$ is learnable, the loss function in the whole model could be depicted as:
@@
Here we keep $f(x)$ frozen and train the side model $g(x)$ and classifier $\sigma$. Unlike a fixed classifier $\sigma^{\prime}$ for text similarity on the original model, our linear-layer-based classifier $\sigma$ can determine if a predicted word exists in the knowledge base of the base language model. This approach reduces inference times for highly repetitive word sequences, storing more general information in the cloud and generating user-specific features on the edge device. Compared to EdgeFormer, our model’s main bottleneck lies in the communication latency[31] between the edge and server. This can be managed by controlling communication frequency and times for an optimal value.
EdgeFormer[30] scales down the original model by factors like four or eight, sharing parameters to achieve a smaller model with similar capabilities and fewer FLOPs. However, this design demands greater computational and proficiency requirements for edge devices. While our methods set the edge model small, its prediction speed relative to total latency is quite small. In this case, the edge devices computational ability would not be a bottleneck for predicting the results on-devices.
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity from back propagate. For prediction, if our transformer has $L$ layers, Parallel adapter[13] or LoRA methods would cause a large cost on communication latency. Specifically, the latency including three parts. $T_{pretrained}$ indicates the original pretrained model computing latency. $T_{on-devices}$ indicates the edge devices model computing latency, it could be Adapter part, LoRA part or LST part for on-devices inference. What’s more, $T_{net}$ is the network latency when transmitting data between cloud and devices. Then, the total latency could be evaluated as:
@@
Where $\mathcal{M}$ is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. LST effectively merges the process in scenarios where edge-cloud connectivity is involved.
@@ 4.2 Implement Details
5 Experiments
Table 1: The Rouge-L scores on LST model, one-shot instruct and SPA model on various datasets.
Model XSum CNN-DM CoQA SciQ
LLaMA-7B + one-shot 17.32 22.36 15.32 17.21
LLaMA-7B + LST 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 35.52 39.22 37.30 25.38
@@ 5.1 Latency on Prediction
LLaMA-7B + S1 1.0 36.24 40.55 38.81 25.91
LLaMA-7B + S2 0.68 30.29 34.19 32.11 23.71
LLaMA-7B + S3 1.0 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 0.62 35.52 39.22 37.30 25.38
Table 4: Different training data sizes on datasets.
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 10000 20000 4000 5000
$\text{SPA}_{\text{medium}}$ 100000 150000 40000 10000
$\text{SPA}_{\text{full}}$ 227000 312000 127000 13679
The network latency for the SPA model is superior compared with other methods. It’s noticeable that SPA exhibits favorable characteristics. It can employ the classifier to choose between a pre-trained model or a model with an integrated adapter for predictions. Due to the LST structure employed on-devices, it offers a more independent architecture compared to methods like LoRA and Adapters.
@@ 5.2 Impact of Side Model
We could find that while the origin cloud model and on-devices model generate different results, the classifier $\sigma$ would make a option that choose a more ideal information from the two alternations. When it comes to a same result, the classifier would choose to pick the cloud one for less network communication cost.
5.3 Analysis on Training Data Sizes
@@
We trained the models using the same network architecture and hyperparameters, with a consistent training epoch of 15 rounds. Text generation was performed using beam search.
@@
Additionally, we recorded the proportion of adapter usage by the model, as shown in Table 6. As data sizes increases, the percentage of adapter usage decreases, with a higher probability of selecting the original model’s output. This further validates that increasing data sizes enhances the model’s inherent language understanding and generation capabilities.
Increasing training data sizes significantly improves downstream task performance. However, there is a diminishing marginal effect with increased data sizes. Considering computational resources, a medium data size provides the best performance-to-cost ratio. More data reduces reliance on adapters, improving the model’s language understanding capabilities.
55551010101015151515202020202424242428282828323232323636363640404040Num of TrainningRouge-LXSumLSTParallel LST
55551010101015151515202020202525252529292929333333333737373741414141Num of TrainningRouge-LCNN-DMLSTParallel LST
Figure 3: Model Rouge-L scores comparison on parallel/sequential side model.
5.4 Different Setting on Side Plugin
SPA outperforms both LST and the pre-trained model, showcasing better expressive capabilities. In order to identify the source of this enhanced performance, we conducted ablation experiments.
We present three configurations in our study. The three configurations we marks from S1 to S3. In the first setup, referred to as "Full Integration, " the edge adapter directly interfaces with the original model. This arrangement allows the edge adapter to effectively learn and correct errors within the original model. In addition, the setup of S1 can be viewed as a parallelized LST setup, which has the advantage that he may perform better on generative tasks than the original serial option of LST.
Table 7: The Side Model Effect evaluation.
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
In the second setup, a choice is made between the edge-adapted model and the original model. The classifier utilizes the original model’s output to make decisions between the results of the original model and the edge-adapted model. In this case, the edge on-devices model operates with a higher degree of independence from the original model.
The final setup involves the existing LST model, which can be viewed as relying entirely on the edge adapter for result generation. By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve. The capabilities demonstrated by different experiments are summarized in Table 3.
The improvement of model performance in downstream tasks primarily relies on the adjustments made by the adapter to capture the incremental differences. As illustrate on Table 3, While the introduction of the classifier in the SPA model leads to a slight performance drop compared to the scenario where adapters are integrated in all cases, the performance reduction is relatively minimal. However, the improvement in inference speed is substantial. Considering our requirement for quick responses at the edge, such a classifier aligns better with our vision and needs. Simultaneously, for more fundamental information, predictions are made using the original model. The classifier significantly reduces the latency caused by continuous calls to the edge-side model during predictions, ensuring that the model learns both user and task features effectively.
We evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L metrics.
5.5 Case Study
@@ 6 Conclusions
This paper introduces a model suitable for cloud-base and on-devices collaboration. This model leverages the capabilities of large models for sequence-to-sequence tasks and utilizes user-specific features from the edge model to guide the overall large model towards our desired objectives. Through a classifier, we can select the most suitable generated sentences. Our approach presents a new paradigm in addressing issues related to edge model deployment, providing valuable insights for future work. We believe that this approach is groundbreaking and can serve as a valuable reference for future research.

### v5

@@ 1 Introduction
Figure 1: Three methods for model deployment are based on different combinations of cloud and device setups.
@@
  • SPA has been thoroughly evaluated across multiple datasets. Experiments show that our framework performs well on-devices. Additionally, SPA can achieve results close to the original model with lower transmission latency.
@@ 2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by [12], incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA [13] employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning [18, 19] involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
@@ 2.2 On-device ML optimization
Given that a considerable portion of these methods can only be applied to models with small model, they may not produce substantial optimizations in the context of large models[36]. Computational and storage resources on devices may not be sufficient to accommodate LLMs. Therefore, we propose a framework for hybrid cloud server and on-device sequence-to-sequence model. The framework aims to meet the needs of certain downstream tasks and user characteristics by combining the capabilities of the cloud and edge.
3 Methodology
Considering the edge computing and storage capabilities, as well as the need to learn user characteristics and contextual information related to text generation, we propose SPA, a lightweight architecture for fast on-devices inference on the constraints of strict on-devices computation and memory constraints. With this design, we can evaluate its support for downstream tasks and compare it with the results of the original edge adapter.
3.1 Side Model Design
Collaborative learning and training between edge and cloud models currently faces the following major bottlenecks[37]: While all models are deployed on edge devices, the computation required for the models involves extensive floating-point operations, and the memory needed for model training significantly increases[38]. These issues make it tough to directly use LLMs for inference on edge devices. On the other hand, when data is placed in the cloud, although general computing power is ensured, some personalized features cannot be efficiently transferred from the edge to the cloud for effective learning. These bottlenecks highlight the urgent need to design a model that combines both edge and cloud capabilities to better adapt to the respective requirements.
We consider placing some learnable parameters on the side to enable rapid personalization of information learning settings. In general, SPA involves using a classifier to select the output results generated by the original LLM or the side block. The classifier chooses the path that aligns better with the desired features, thereby optimizing the quality of text generation. Additionally, the classifier can decide whether to combine the edge model with the base model stored in the cloud. In this way, the edge model only needs to learn the feature differences relevant to the current task, while the general information is provided by the base model, reducing the resource latency costs on edge devices.
@@ 3.1 Side Model Design
$\mathcal{H}_{i}(x)=f_{i}(x)+\sigma_{i}g_{i}(x)$ (1)
We employed a linear layer of $H\times 2$ as our classifier. The advantage of this design lies in our ability to use the Transformer’s output as the hidden state. This hidden state is then softmaxed to a range between $[0,1]$, thereby obtaining a specific distribution for classification choices $\sigma_{i}=\arg\max{\textbf{M}_{c}f_{i}(x)}$.
Through such a linear mapping, the model can learn whether predicting a certain token is better accomplished by selecting the existing model’s output. This approach typically leverages previously predicted information like: Los Angeles, Which Los is predicted in previous token. Alternatively, the model can employ a side-end incremental model that relies more on user-specific feature information. This helps in determining the optimal choice. Since $\sigma_{i}$ is learnable, the loss function for the whole model can be depicted as follows:
Figure 2: Cloud and Device collaboration model SPA. This framework utilizes a classifier to provide optional downstream task computation choices for edge devices, thereby reducing computational pressure on edge devices and improving the overall performance of specialized device models.
@@
Here, we keep $f(x)$ frozen and unchanged, while training the side model $g(x)$ and the classifier $\sigma$. Unlike the fixed classifier $\sigma^{\prime}$ used for text similarity in the original model, our classifier $\sigma$ based on a linear layer can determine whether the predicted word exists in the knowledge base of the underlying language model. This approach reduces the inference time for highly repetitive word sequences, stores more general information in the cloud, and generates user-specific features on edge devices. Compared to EdgeFormer, the main bottleneck of our model is the communication latency between the edge device and the server [17]. This can be optimized by controlling the frequency and timing of communication. EdgeFormer [32] achieves a smaller model with similar functionality but fewer FLOPs by reducing the original model’s size (e.g., by four or eight times) and sharing parameters. However, this design requires higher computational power and specialization for edge devices. While our approach results in a smaller edge model, its prediction speed relative to the total latency is very fast. In this case, the computational power of the edge device will not become a bottleneck for the prediction results.
3.2 Communications Latency
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity through backpropagation. For prediction, if our transformer has $L$ layers, the Parallel adapter or LoRA method results in higher communication latency costs. Specifically, the latency includes three parts: $T_{pretrained}$ represents the computation latency of the original pre-trained model, $T_{on-devices}$ represents the computation latency of the edge device model, which could be the Adapter part, LoRA part, or LST part used for edge device inference. Additionally, $T_{net}$ is the network latency when transmitting data between the cloud and devices. Then, the total latency can be evaluated as:
@@ 3.2 Communications Latency
Where $\mathcal{M}$ is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. SPA effectively merges the process in scenarios where edge-cloud connectivity is involved.
@@ 4.1 Datasets
4.2 Baselines
We compare our approach with various efficient parameter tuning methods, including three baseline: LoRA[13], Adapters[12], and LST[21]. These methods are widely used in different downstream tasks and have achieved outstanding results.
4.3 Models
We utilize Llama[3], an open-source large language model, and employed the PEFT method to evaluate task inference latency, network latency, and task performance. We conducted our analysis using the Llama-7B model. This open-source large language model demonstrates very strong capabilities in general domains.
@@ 5.2 Latency on Prediction
LLaMA-7B + Fully Integration 1.0 36.24 40.55 38.81 25.91
LLaMA-7B + Fully Separation 0.68 30.29 34.19 32.11 23.71
LLaMA-7B + LST 1.0 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 0.64 36.52 39.22 37.30 25.38
The SPA model outperforms other methods in terms of network latency. Notably, SPA has many excellent features. It can make predictions by choosing between the pre-trained model and the integrated adapter model through a classifier. Due to the use of the LST structure, SPA’s architecture is more independent on devices, giving it an advantage over methods like LoRA and Adapters.
@@ 5.3 Impact of Side Model
We could find that while the origin cloud model and on-devices model generate different results, the classifier $\sigma_{i}$ would make a option that choose a more ideal information from the two alternations. When it comes to a same result, the classifier would choose to pick the cloud one for less network communication cost.
5.4 Training Scales Analysis
@@
We train the models using the same network architecture and hyperparameters, with a consistent training epoch of 15 rounds. Text generation is performed using beam search.
@@
Additionally, we record the proportion of adapter usage by the model, as shown in Table 5. As data sizes increases, the percentage of adapter usage decreases, with a higher probability of selecting the original model’s output. This further validates that increasing data sizes enhances the model’s inherent language understanding and generation capabilities. Increasing the size of training data significantly enhances performance on downstream tasks. However, the marginal benefits decrease as data size continues to grow. Considering computational resources, a medium-sized dataset offers the best performance-to-cost ratio. Additionally, more data reduces the need for adapters, thereby improving the model’s language understanding capabilities.
5.5 Alternatives on Side Plugin
SPA outperforms both LST and the pre-trained model, showcasing better expressive capabilities. we conduct ablation experiments for identifying the source of this enhanced performance.
We present three configurations in our study. In the first setup, referred to as Fully Integration. The edge adapter directly interfaces with the original model. This arrangement allows the edge adapter to effectively learn and correct errors within the original model. In addition, the setup of S1 can be viewed as a parallelized LST setup, which has the advantage that he may perform better on generative tasks than the original serial option of LST.
Figure 3: Performance on XSum datasets between LST and SPA.
Figure 4: Performance on CNN-DM datasets between LST and SPA.
In the second setup, a choice is made between the edge-adapted model and the original model. The classifier utilizes the original model’s output to make decisions between the results of the original model and the edge-adapted model. We denote the setting as Fully Separation In this case, the edge on-devices model operates with a higher degree of independence from the original model.
Table 4: The Side Model Effect evaluation.
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
The final setup involves the original LST framework, which can be viewed as relying entirely on the edge adapter for result generation. By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve. The capabilities demonstrated by different experiments are summarized in Table 2.
The performance improvement in downstream tasks mainly depends on the adapter’s adjustments to capture incremental differences. As shown in Table 2, while introducing the classifier in the SPA model causes a slight performance drop compared to integrating adapters in all scenarios, this reduction is minimal. However, the increase in inference speed is significant. Given our need for quick responses at the edge, this classifier aligns well with our vision and requirements. For more fundamental information, predictions are made using the original model. The classifier greatly reduces latency from continuous calls to the edge-side model during predictions, ensuring effective learning of both user and task features.
Parallel processing. We also evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L metrics.
5.6 Case Study
@@ 6 Conclusions
This paper introduces a model suitable for cloud-base and on-devices collaboration. The model leverages the capabilities of LLMs for sequence-to-sequence tasks and utilizes user-specific features from the edge model to guide the overall large model towards our desired objectives. Through a classifier, we can select the most suitable generated sentences. Our approach presents a new paradigm in addressing issues related to edge model deployment, providing valuable insights for future work. We believe that this approach is groundbreaking and can serve as a valuable reference for future research.
Appendix A Limitation
Our model is a cloud-based and on-device cooperative system, which exhibits a strong inference ability on various downstream tasks. Leveraging the high-parameter nature of the large model, it can achieve excellent end-to-end prediction results. However, our model still has some shortcomings. For instance, its collaborative process inevitably introduces some latency. In scenarios that demand high network resources, our model may not perform optimally for reasoning and training.
Additionally, there may be privacy concerns since our corpus is ultimately processed in the cloud, allowing cloud services to access the corpus content. Nonetheless, we protect privacy to a certain extent by keeping the model-related parameters locally. We only need these local parameters to perform inference based on user features. These issues warrant further investigation and improvement in future work.